(Prepared from a situation suggested by Professor John W. Hardy.) Lone Star Meat Packers is a major processor of beef and other meat products. The company has a large amount of T-bone steak on hand, and it is trying to decide whether to sell the T-bone steaks as they are initially cut or to process them further into filet mignon and the New York cut.

If the T-bone steaks are sold as initially cut, the company figures that a 1-pound T-bone steak would yield the following profit:

Selling price ($7.95 per pound) $ 7.95
Less joint costs incurred up to the split-off point where
T-bone steak can be identified as a separate product
3.80
Profit per pound $ 4.15

If the company were to further process the T-bone steaks, then cutting one side of a T-bone steak provides the filet mignon and cutting the other side provides the New York cut. One 16-ounce T-bone steak cut in this way will yield one 6-ounce filet mignon and one 8-ounce New York cut; the remaining ounces are waste. It costs $0.55 to further process one T-bone steak into the filet mignon and New York cuts. The filet mignon can be sold for $12.00 per pound, and the New York cut can be sold for $8.80 per pound.

Required:

1. What is the financial advantage (disadvantage) of further processing one T-bone steak into filet mignon and New York cut steaks?

2. Would you recommend that the T-bone steaks be sold as initially cut or processed further?
Answer #1
Concepts and reason

Profit margin: It can be stated as the amount of revenue in excess of the expenditure. It is used in evaluating the profitability of the operations of the business.

Fundamentals

Sales Revenue: The income or amount which is received from services provided and sale of goods by the company which also includes the discounts and deductions during the year is called sales revenue.

Inventory: It is goods of the company which can be in forms of raw material, work in progress and finished goods. The company held these goods for sale in the ordinary course of business it shows the capability of the company to manufactured goods in what quantity it is showed in the asset side of the balance sheet.

Sunk Cost: It is the cost which does not affect the investment decision. The sunk cost always occurs and it is not dependent on any event.

Joint cost: This is the cost involved in the joint process. The joint cost involves the cost of direct material, direct labor and overhead cost involved in the joint production process.

Cost and Expenses: Cost is the amount expended to any particular product or amount related to any particular product. Whereas, expenses means the cost incurred in order to earn income. Prepaid Expense and Outstanding Expense are its two types. It is shown on the debit side of profit and loss Account

1.

Compute the financial advantage (disadvantage) of further processing of T-bone into filet mignon and New York cut steaks using the equation as shown below:

Financialadvantage=(TotalsalesfromfurtherprocessingSalerevenuelostofoneTboneCostoffurtherprocessing)=$8.90$7.95$0.55=$0.40\begin{array}{c}\\{\rm{Financial advantage}} = \left( \begin{array}{l}\\{\rm{Total sales from further processing}} - \\\\{\rm{Sale revenue lost of one T - bone}} - \\\\{\rm{Cost of further processing}}\\\end{array} \right)\\\\ = \$ 8.90 - \$ 7.95 - \$ 0.55\\\\ = \$ 0.40\\\end{array}

Hence, the financial advantage (disadvantage) of further processing of T-bone into filet mignon and New York cut steaks is $0.40.

Working Notes:

Compute the total sales from further processing using the equation as shown below:

Salesfromfurtherprocessing=(OneFiletedMignon+NewYorkCut)=$4.50+$4.40=$8.90\begin{array}{c}\\{\rm{Sales from further processing}} = \left( \begin{array}{l}\\{\rm{One Fileted Mignon}} + \\\\{\rm{New York Cut}}\\\end{array} \right)\\\\ = \$ 4.50 + \$ 4.40\\\\ = \$ 8.90\\\end{array}

Hence, the total sale from further processing is $8.90.

Compute the Sales revenue from one fileted mignon after further processing using the equation as shown below:

OneFiletedMigon=(Sellingpriceperfiletmignon×YeildperounceSizeofoneTbonesteak)=$12×6ounce16ounce=$4.50\begin{array}{c}\\{\rm{One Fileted Migon}} = \left( {\frac{{{\rm{Selling price per filet mignon}} \times {\rm{Yeild per ounce}}}}{{{\rm{Size of one T - bone steak}}}}} \right)\\\\ = \frac{{\$ 12 \times 6{\rm{ ounce}}}}{{16{\rm{ ounce}}}}\\\\ = \$ 4.50\\\end{array}

Hence, the sales revenue from one fileted mignon after further processing is $4.50.

Compute the Sales revenue from one New York cut after further processing using the equation as shown below:

NewYorkcut=(SellingpriceperNewYorkcut×YeildperounceSizeofoneTbonesteak)=$8.8×8ounce16ounce=$4.40\begin{array}{c}\\{\rm{New York cut}} = \left( {\frac{{{\rm{Selling price per New York cut}} \times {\rm{Yeild per ounce}}}}{{{\rm{Size of one T - bone steak}}}}} \right)\\\\ = \frac{{\$ 8.8 \times 8{\rm{ ounce}}}}{{16{\rm{ ounce}}}}\\\\ = \$ 4.40\\\end{array}

Hence, the sales revenue from one New York cut after further processing is $4.40.

2.

The T-bone steaks should be processed further as further processing of T-bone steaks would lead to the net advantage of $0.40.

Ans: Part 1

The financial advantage (disadvantage) from further processing is $0.40.